### v3

# DPOD: Domain-Specific Prompt Tuning for Multimodal Fake News Detection

Abstract
The spread of fake news using out-of-context images has become widespread and is a relevant problem in this era of information overload. Such out-of-context fake news may arise across different domains like politics, sports, entertainment, etc. In practical scenarios, an inherent problem of imbalance exists among news articles from such widely varying domains, resulting in a few domains with abundant data, while the rest containing very limited data. Under such circumstances, it is imperative to develop methods which can work in such varying amounts of data setting. In this work, we explore whether out-of-domain data can help to improve out-of-context misinformation detection (termed here as multi-modal fake news detection) of a desired domain, to address this challenging problem. Towards this goal, we propose a novel framework termed DPOD (Domain-specific Prompt-tuning using Out-of-Domain data). First, to compute generalizable features, we modify the Vision-Language Model, CLIP to extract features that helps to align the representations of the images and corresponding text captions of both the in-domain and out-of-domain data in a label-aware manner. Further, we propose a domain-specific prompt learning technique which leverages the training samples of all the available domains based on the extent they can be useful to the desired domain. Extensive experiments on a large-scale benchmark dataset, namely NewsCLIPpings demonstrate that the proposed framework achieves state of-the-art performance, significantly surpassing the existing approaches for this challenging task. Code will be released on acceptance.
Keywords:
Multi-modal fake news detection Prompt tuning Vision-language models
1 Introduction
In this digital world, social media has become the main source of information for an increasing fraction of the population. Thus, it is very important that the news reaching the masses is authentic, since fake news can have serious consequences, like manipulating public opinion, stirring up conflicts, and even impacting financial markets. With the internet connecting billions of people worldwide, fake news can spread at an alarming speed, thus it is equally important to identify it quickly and efficiently. One increasingly common form of fake news is the out-of-context use of images, where a real image is paired with a false caption before dissemination. It is very difficult to detect such fake news, since the image is real, i.e., not manipulated or generated, and it is only recently that researchers have started to address this challenging, but very important problem [2][1].
In a real-world setting, there is an implicit skewness present in data across varying domains. For example, there is an increased dominance of news domains like politics over certain specific topics like education. However, one maybe interested in news from a particular domain, e.g., a doctor may be more interested in healthcare news, whereas a government official may be more interested in political news. Due to the presence of such widely varying data quantities across different domains, it is difficult to correctly predict the veracity of a news sample belonging to a particular domain of interest. However, news articles may often share overlapping traits, with different domains intersecting with each other. Hence, for a target domain of choice, we can leverage information from other domains to address this problem of imbalance.
In this work, we propose a novel framework termed DPOD (Domain-specific Prompt-tuning using Out-of-Domain data) for out-of-context misinformation detection by leveraging both image and caption pairs, which we term here as MFND or Multi-modal Fake News Detection. It is evident that training a model on the samples of a specific domain will yield good performance when evaluated on the same domain. However, this is not practical, since in addition to training different models for different domains, training on domains with less data may be difficult. Towards this goal, we develop a generalized framework, where we consider data from all domains while training. The model is trained such that it is simultaneously tailored to each individual domain, by leveraging information from all the available domains. As an example, for MFND on a particular domain of interest like politics, we effectively utilize out-of-domain data from healthcare, international relations, etc. to improve its performance. The DPOD framework first utilizes label-aware alignment of the available image-text pairs of all the domains to learn generalizable features, irrespective of the domain of interest. In our work, we harness the capabilities of a large Vision Language Model (CLIP) [18] to comprehend interwoven text and image features. Next, a semantic domain vector is computed for each of the training data domains, which incorporates the similarity of a particular domain with all the other available training domains. Further, this is utilized to learn a domain-specific prompt, along with a set of generic learnable prompts, which is finally used for predicting whether a given image-caption pair is real or fake. During inference, the model specialized to the test data can be utilized, even when the domain label of the test data is unknown. Extensive experiments on a large-scale benchmark dataset, namely NewsCLIPpings [15] demonstrate the effectiveness of DPOD over state-of-the-art approaches. The contributions of this work can be summarized as follows:
  1. 1.
    We propose a novel framework, termed DPOD, which effectively utilizes out-of-domain data for improving the performance of in-domain data, hence addressing the issue of varying amounts of data across different domains.
  2. 2.
    We utilize a label-aware loss to learn a generalizable model, and propose a novel framework to learn domain-specific prompts using out-of-domain data.
  3. 3.
    This helps in quick model deployment for the target domain, even when it contains limited domain-specific annotated data, thus being effective for scarcely annotated data scenarios.
  4. 4.
    Extensive experiments show that the proposed DPOD outperforms the existing approaches in different settings, thus achieving the new state-of-the-art.
Now, we briefly describe the related work, followed by the proposed framework and the experimental evaluation.
2 Related Work
Here, we briefly describe the related work in MFND, self-supervised learning and prompt tuning.
Multi-modal Fake News Detection: Recently due to the rapid spread of fake news using out-of-context real images and texts/captions, researchers have started to address this socially relevant important problem. [15] proposed a novel dataset named NewsCLIPpings, derived from the Visual News dataset [14] and suggested a baseline solution using CLIP and VisualBERT feature extractor. The SAFE framework [29] utilises Text-CNN and VGG [20] model for extracting features from the texts and their associated images which are then used to predict the fake news. Similarly, [22] use BERT [5] as text encoder to extract textual features and use VGG-19 as image encoder to extract visual features and then finally fuse them together for classification. Most of these existing methods fail to perform well on unseen events. [23] trained an event discriminator network concurrently with multi-modal fake news detector to eliminate event-specific characteristics while retaining the common features among different events. The work in [19] jointly preserves domain-specific and cross-domain knowledge in news records to detect fake news from different domains by using combination of domain-specific and cross-domain features. In [24], meta-learning and neural process methods are integrated to achieve high performance even on events with limited labeled data.
The emergence of large scale vision-language models (VLMs) which learns information from both images and texts, has proved to be immensely useful for computer vision tasks. Models like CLIP [18] and ALIGN [11] have been trained in a contrastive manner on large scale datasets, such that they learn rich representations between images and texts. [30] used CLIP along with image encoder (ResNet) and text encoder (BERT) to extract the multimodal features and used attention over different modalities to get the final features for classification. Recently, [17] proposed a novel DT-Transformer framework for MFND using CLIP encoders and a Transformer-based detector instead of using a binary classifier. Some recent works also use social media related information to detect the fake news [19][1][25]. [19] proposed a framework which uses image, text and propagation graph for MFND and [25] suggested a graph model based on CLIP embeddings. A Consistency-Checking Network is proposed in [1] to measure the consistency of the image and text pairs along with the evidences collected from the internet.
Self-Supervised Learning: Self-supervised learning leverages the inherent structure of data to generate its own supervisory signals without relying on large labeled datasets. [4] have showed that the combination of data augmentations has a crucial impact on useful predictive tasks. Supervised contrastive loss [13] leverages label information for self-supervised representation learning. It pulls together clusters of points belonging to the same class while pushing apart clusters of samples from different classes. Recently, self-supervision has been used successfully for out-of-context misinformation detection using images and text. In [2], self-supervised learning is used to align individual objects in an image with the corresponding textual claims, where, given two captions, the goal is to detect whether an image is being used out-of-context. The state-of-the-art method  [16] propose a Self-Supervised Distilled Learner, which uses self-supervised learning to detect fake news, where they propose a Teacher network to guide a Student network in mimicking similar decision patterns. Though we also utilize self-supervision, our proposed DPOD framework is much simpler, yet more effective, and does not rely on teacher-student model.
Vision-Language Models and Prompt Tuning: Since, adapting large VLM’s like CLIP to downstream tasks may prove to be difficult, methods like prompt learning have come up as a popular research direction. The text instructions given to the text encoder, are called “Prompts”. Prompt learning is an alternative to handcrafted prompts, where the prompt tokens are learned during the finetuning stage. Recently, many works have proposed prompt learning techniques such as CoOp [28], CoCoOp [27], Maple [12], etc., which increases the generalizability of the VLMs. Some works specifically address the domain adaptation problem, for example, [26] proposes a domain-invariant prompt tuning, which can generalize to unseen domains. [9] uses prompt tuning in a unimodal context (in language models), whereas, [3] uses only visual prompt tuning. The works in [8, 21, 6] learns prompts specifically for each domain. To the best of our knowledge, ours is the first work to explore prompt-tuning for the challenging MFND task.
3 Problem Definition and Model Overview
The goal of the proposed framework is to leverage out-of-domain training samples when testing on a certain target domain. Let the train dataset be denoted as $\mathcal{S}_{train}=\{{D}_{1},{D}_{2},...,{D}_{n}\}$. ${D}_{i}$ denotes each domain and can be written as, ${D}_{i}=\{({I}_{k},{T}_{k},{y}_{k})\},k=1,2,...,{n}_{{D}_{i}}$, where ${n}_{{D}_{i}}$ is the number of samples in domain ${D}_{i}$. ${I}_{k}$ and ${T}_{k}$ denote the $k^{th}$ image and text pair, and ${y}_{k}$ denotes the ground truth, i.e., fake or true news. The news samples can originate from various sources (or domains) such as politics, healthcare, sports, etc. We assume that we know the domain of the training examples, though we show empirically that the proposed framework can inherently handle domain label inconsistencies that is expected from crowd-sourced data annotation. We assume that an user is interested in a specific domain $D\in\{{D}_{1},{D}_{2},...,{D}_{n}\}$ during testing, though our model training happens only once irrespective of the desired target domain.
During testing, we want to predict whether a given image-text pair is fake or real. First, we briefly describe the base model used.
Base Architecture. Recently, vision language foundation models like CLIP [18], Align [11] etc., have been immensely successful in relating images and the corresponding text and have also been used for the MFND task [16][1]. CLIP is trained using natural language, which being expressive, can supervise a much broader set of visual concepts. Following the recent literature, we also use a modified CLIP architecture for our work. The broader structure of our proposed DPOD framework can be divided into three stages as follows (Fig. 1):
Stage 1: Label-Aware Alignment of Multi-modal data: In this stage, we train all the layers of the original CLIP model in an end-to-end manner using the label-aware alignment loss on all the training data, irrespective of their domains, to obtain the aligned clip, termed here as A-CLIP.
Stage 2: Creating Semantic Domain Vectors. Here, we freeze the already learnt image and text encoders of the A-CLIP model. Using the domain-wise joint image-text embeddings from these frozen encoders, we create semantic domain vectors for each of the training domains. Each index of this vector denotes its similarity with the joint embedding of all the other domains.
Stage 3: Domain-Specific Prompt tuning with Out-of-Domain data: With the above settings, we next propose a prompting strategy during training. Following previous works [28][27], we append a prefix prompt to the text caption. However, a fixed prompt like “A photo of” [18], may not be the best prefix for all the captions. Hence, we make the prefix prompt tokens learnable as $\{{V}_{1},{V}_{2},{V}_{3}\}$. Additionally, we pass the semantic domain vector though a linear layer $\mathcal{F}_{\theta}$, and get a domain-specific prompt token, which is appended to the prefix prompt tokens. The complete text captions appended with the learnable prompts, and the corresponding images are passed though the frozen encoders to create the final joint embedding, which is subsequently classified into fake or real news.
4 Proposed DPOD Framework
Here, we describe each of the three stages of the DPOD framework presented in Fig. 1 in detail.
4.1 Stage 1: Label-Aware Alignment
Figure 1: Illustration of the different stages of the proposed DPOD approach. In Stage 1, we learn the image and text encoders via Label-Aware Alignment Loss. Then this model is used to obtain semantic domain vectors in Stage 2. This, in turn is finally used to learn domain-specific and generic prompts with out-of-domain samples to predict the veracity of the news in Stage 3.
Since our goal is to predict the veracity of news (image-caption pair) in a particular domain ($D$, say politics), the ideal scenario is when vast amounts of annotated data for domain $D$ are available. In real scenarios, getting large amounts of annotated data from each domain is challenging. Our hypothesis is that there might be overlapping concepts/objects in the data from different domains, like healthcare, entertainment, which can be utilized to mitigate the adverse effect of less amount of training data of the desired domain.
Using self-supervision as a warm-up stage for model training has been used successfully in previous works like [2, 16]. In contrast to these approaches, where both the training and test sets contain data from all (but same) domains, we explore whether self-supervision from out-of-domain data can benefit the performance of in-domain data for MFND task. Here, the original CLIP model is trained to learn relations between texts and images from all the domains, but in a label-aware manner. Specifically, we use a multi-modal contrastive loss, such that the backbone model learns to bring closer the augmentations of the image and the corresponding text when the news is real and can separate them when the news is fake.
For label-aware alignment, initially, four augmentations, namely, color jitter, random crop, horizontal flipping and normalisation are generated per image. Then the CLIP embeddings of the original image, text and the image augmentations are generated. Suppose an image and the corresponding text be denoted as $I_{k}$ and $T_{k}$ respectively and the image augmentations be denoted by $I_{ak},a=1,\ldots,4$. We denote the corresponding embeddings as $\{\hat{I}_{k},\hat{T}_{k},\hat{I}_{ak}\}$. For the consistent (real) image-text pairs, we want to bring the embeddings of the image ($\hat{I}_{k}$), its augmentations ($\hat{I}_{ak},a=1,\ldots,4$) and the corresponding text ($\hat{T}_{k}$) close to one another. Thus for the positive pair, we have two combinations, namely image-image pair and image-text pair. The corresponding label-aware alignment losses for the $i^{th}$ image augmentation of $k^{th}$ input sample are calculated as:
[…]
where $\hat{I}_{ik}$ and $\hat{I}_{jk},j\neq i$ denotes the $i^{th}$ and $j^{th}$ augmentations of the image $I_{k}$. $\tau$ is the temperature parameter that controls the smoothness of the probability distribution, $N$ is the number of training image-text pairs and $sim$ represents the cosine similarity. We use $\tau=0.05$ for all our experiments. The denominator $\mathcal{D}_{true}$ represents the combination of all the negative pairs and is given as
$\mathcal{D}_{\text{true}}=\sum_{\begin{subarray}{c}t=1\\ t\neq k\end{subarray}}^{N}\left(\sum_{a=1}^{4}\exp(\text{sim}(\hat{I}_{ik},\hat{I}_{at}))+\exp(\text{sim}(\hat{I}_{ik},\hat{T}_{t}))+\exp(\text{sim}(\hat{T}_{k},\hat{T}_{t}))\right)$
The total loss corresponding to the real or true news is obtained by adding the losses in eq. (1):
[…]
For the inconsistent (fake) image-text pair, the label-aware alignment of the images with their augmentations is computed in the same manner as in the true case. Since the image-text pair is inconsistent, they are pushed apart unlike in the true case. Thus the final loss for the fake pair is:
[…]
Here, the denominator is given as
$\mathcal{D}_{\text{fake}}=\sum_{\begin{subarray}{c}t=1\\ t\neq k\end{subarray}}^{N}\left(\sum_{a=1}^{4}\exp(\text{sim}(\hat{I}_{ik},\hat{I}_{at}))+\exp(\text{sim}(\hat{T}_{k},\hat{T}_{t}))\right.$
$+\left.\exp(\text{sim}(\hat{I}_{ik},\hat{T}_{t}))\right)+\exp(\text{sim}(\hat{I}_{ik},\hat{T}_{k}))$
The final label-aware alignment loss is computed by taking a weighted sum of the two losses corresponding to real and fake data using the eq. (4) below
[…]
The parameter $\beta$ is used to weigh the two terms. This differential treatment of fake and real samples in both stages helps to better train the model and reduce confusion.
4.2 Stage 2: Computing Semantic Domain Vectors
In this stage, we create semantic domain vectors which represent how similar each domain is to all the domains in the training dataset. If the number of domains in the training dataset $\mathcal{S}_{train}$ is $n$, we create $n$ semantic domain vectors each having $n$ indices. Let us take the “politics” domain in the training dataset as an example. To compute the semantic domain vector for “politics”, we pass all its images and corresponding text captions, i.e., $\{({I}_{k},{T}_{k})\},k=1,2,...,{n}_{{D}_{i}}$ through the frozen encoders of the A-CLIP trained in Stage 1 to get $\{({\hat{I}}_{k},{\hat{T}}_{k})\}$. Here, ${n}_{{D}_{i}}$ denotes the number of training samples in the politics domain. We take a Hadamard product of the image and text embedding $\{({\hat{I}}_{k},{\hat{T}}_{k})\}$ to get ${n}_{{D}_{i}}$ joint embedding. Then, we compute the mean across all the sample joint embedding to get the final mean joint embedding $\mathcal{J}_{{D}_{i}}$, particular to a domain ${D}_{i}$ (here, “politics”) as shown below:
$\mathcal{J}_{{D}_{i}}=\frac{\sum_{k=1}^{{n}_{{D}_{i}}}\mathcal{J}_{k}}{{n}_{{D}_{i}}},where,\mathcal{J}_{k}=\hat{I}_{k}\odot\hat{T}_{k}$ (5)
This is illustrated in Fig. 1(a) Finally, to create the semantic domain vector ${w}_{{D}_{i}}$, we take the cosine similarity between this joint embedding $\mathcal{J}_{{D}_{i}}$ and those of the other training domains $\mathcal{J}_{{D}_{k}},k=1,2,...,n$. Hence, the $j^{th}$ index of the ${w}_{{D}_{i}}$ vector will be the cosine similarity of $\mathcal{J}_{{D}_{i}}$ with $\mathcal{J}_{{D}_{j}}$, ($j\in k$) as shown in eq. (6).
${{w}_{{D}_{i}}(j)}=\frac{{\mathcal{J}_{{D}_{i}}\cdot\mathcal{J}_{{D}_{j}}}}{\|\mathcal{J}_{{D}_{i}}\|\|\mathcal{J}_{{D}_{j}}\|}$ (6)
We term these as semantic domain vectors, since they can implicitly capture the semantics of the domains, i.e. the domain vectors of similar domains are closer compared to those of unrelated domains. This is illustrated in detail later.
4.3 Stage 3: Domain-Specific Prompt-tuning with Out-of-Domain data
In this stage, we utilize the semantic domain vectors to learn domain-specific prompt embedding that help to incorporate domain-specific information in the model while classifying a news item as true or fake. We initialize a prompt prefix with the standard text “A photo of” [18]. However, we make this prefix learnable (similar to [28]), since it may not be the best choice for our application. We denote these generic learnable prompts as $\{{V}_{1},{V}_{2},{V}_{3}\}$. To incorporate the domain information, we pass the corresponding semantic domain vector ${w}_{{D}_{i}}$ (obtained in the previous stage) through a linear layer, denoted by $\mathcal{F}_{\theta}$. This layer $\mathcal{F}_{\theta}$ projects the ${n}_{{D}_{i}}$-dimensional domain vector to a 512-dimensional domain prompt token $\mathcal{V}_{{D}_{i}}$, which is then appended to the prompts $\{{V}_{1},{V}_{2},{V}_{3}\}$. This forms the whole learnable prompt structure, where $\{{V}_{1},{V}_{2},{V}_{3}\}$ is trained in a domain-agnostic manner, whereas, $\mathcal{V}_{{D}_{i}}$ is trained by conditioning on the domain of the particular training sample. This entire prompt is then prepended to the text caption ${T}$ and passed through the frozen A-CLIP text encoder to get a text embedding $\mathcal{\hat{T}}$ as shown in eq. (7).
$\mathcal{\hat{T}}=\{\theta_{1},\theta_{2},\theta_{3},\mathcal{F}_{\theta}({w}_{D}),T\}$ (7)
[…]
The Classifier Network consists of two fully connected layers which project $\mathcal{\hat{J}}$ to $\mathcal{\hat{X}}$. Inspired by [7], we also introduce a residual connection from $\mathcal{\hat{J}}$ to $\mathcal{\hat{X}}$, the output of which finally passes through two fully connected layers to give a final prediction $\hat{y}$. The whole network is trained in an end-to-end manner using a Binary Cross-Entropy loss as shown in eq. (8).
$\mathcal{L(\theta})=-{(y\log(\hat{y})+(1-y)\log(1-\hat{y}))}$ (8)
Here, $\theta$ denotes all the learnable parameters in the network. The ground-truth ${y}=1$, if the sample is fake, and ${y}=0$, if it is a true news.
Inference. During inference, we consider two scenarios: one where we have the domain information for the test image-text pair, and another where the target domain is unknown. For the first case, we feed its corresponding semantic domain vector through the already trained $\mathcal{F}_{\theta}$ network to get the target specific domain prompt ${V}_{D}$. For the second case, we find the most similar domain of the image-text sample by comparing its joint embedding with all the domain-specific mean joint embeddings, since the domain is unknown. Specifically, we find the domain $D_{t}$ of this test pair as follows: $D_{t}=\operatorname*{argmax}_{D}(\mathcal{J}_{t}\cdot\mathcal{J}_{D_{i}}),i=1,2,...,n$, where, $\mathcal{J}_{t}$ denotes the joint embedding of the test sample, and $\mathcal{J}_{D_{i}}$ denotes the domain-wise mean joint embeddings. Next, the text caption (of test data) is appended with the learnt prompt vectors to get $\{{V}_{1},{V}_{2},{V}_{3},{V}_{D}\}$. The final text and image embedding obtained using the frozen A-CLIP encoders are then used to compute the joint embedding. The Classifier Network gives the final prediction ${\hat{y}}$, which is thresholded to give a binary prediction as True or Fake.
Difference with other prompt-learning techniques. As discussed in the Related Work section, there has been a recent few works on domain-based prompt tuning [8, 21]. But there are few important differences between these techniques and the proposed one:
1) In the existing works, “domain” or “style” refers to changes in image features. Contrary to this, in our case, domains (such as “politics”, “law-crime”, etc.) do not have significant change in image features (since, images from multiple domains will contain similar objects like humans, etc.), but are subjective to human interpretation of both the images and the texts jointly.
2) In existing works, the text only contains the class name as opposed to a sentence describing the image as in ours.
3) We propose a novel strategy of creating semantic domain vectors, which leverages similarity from out-of-domain data, and utilize it to create domain-specific learnable prompts. To the best of our knowledge, this type of domain-prompting strategy has not been explored earlier.
Table 1: Comparison of the performances of our proposed method against existing approaches on the NewsCLIPpings dataset where we train and test on all the domains.
Methods Backbone Accuracy (%) DT-Transformer [17] CLIP ViT-B/32 65.67 MNSL [25] CLIP 68.20 DCD [10] CLIP 65.40 NewsCLIPpings [15] CLIP ViT-B/32 60.23 CLIP-FT [1] CLIP ViT-B/16 64.79 CLIP ViT-B/32 65.10 SSDL [16] CLIP ViT-B/16 65.00 CLIP RN-50 71.00 DPOD (Ours) CLIP ViT-B/16 70.49 DPOD (Ours) CLIP ViT-B/32 70.81 CLIP RN-50 74.43
5 Experimental Evaluation
We evaluate the effectiveness of the proposed framework on the standard large-scale benchmark, namely NewsCLIPpings dataset [15], which is created from the Visual News dataset [14]. It contains news articles from sources like: The Guardian, BBC, USA Today, and The Washington Post. It has a total of $71,072$ train, $7,024$ validation and $7,264$ test examples. It contains news articles belonging to a total of 54 domains with varied amounts of training data across each of the domains for example, politics (1396 samples), law-crime (4348 samples), sports (2786 samples) etc., effectively relecting the imbalance problem.
Implementation Details. CLIP “ViT-B/32”[18] is the backbone of all our experiments. We use accuracy as the evaluation metric as in [1, 16], as the data has equal distribution of true and fake news. Our codes are implemented in PyTorch and uses a single NVIDIA GeForce RTX 3080 Ti GPU. ADAM optimizer with weight decay and stability constant is used for the training. All the results are reported using the same set of parameters, e.g., $\beta=1.5$, $lr=1e-4$ and $batchsize=64$ across all dataset splits. The values of hyperparameters are obtained during the validation process against the constant validation dataset containing all-domain data. These values were then kept fixed for the rest of the experiments. Stage 1 is run for 40 epochs, which remains fixed even if the desired domain changes. We initialize our domain-agnostic prompts $\{V_{1},V_{2},V_{3}\}$ required in Stage 3, as “A photo of” which is then learnt through training. After obtaining the output probability at the end of Stage 3, we use Sigmoid function to convert the output into a probability score $\hat{y}$ between 0 to 1. We train Stage 3 for 30 epochs. At the time of inference, the data having output greater than or equal to $0.5$ is assigned the label 1, i.e., Fake and the others are assigned label 0, i.e., True. Now, we describe the experiments conducted to evaluate the effectiveness of the proposed DPOD framework. We would specifically like to answer the following research questions:
  1. 1.
    How does the model perform for the standard setting, where all the domain data is used for both training and testing?
  2. 2.
    How well does the proposed model perform for a desired target domain under the imbalance scenario?
  3. 3.
    Can our model generalize to completely unseen target domains?
  4. 4.
    Can the model handle inconsistencies in domain labels and how well does it understand about the domains?
  5. 5.
    Are all the components in the model useful?
Table 2: Comparison when the trained model is tested on target datasets having varying number of training samples (in parenthesis). Top (bottom) section contains results on domains having highest (lowest) number of samples.
Model Maximum and Least Represented Domains
Sport(3518)
Law
Crime(4348)
Internatn.
Relations(5832)
World(5238)
Arts and
Culture(13736)
CLIP-FT[1] 68.32 59.92 68.76 62.11 68.52
CoOp[28] 69.68 62.87 69.13 64.37 69.61
DPOD (Ours) 71.01 67.82 69.47 65.24 73.49
Fashion(284)
TV
Radio(470)
Education(376) Money(224) Travel(526)
CLIP-FT[1] 56.68 59.90 63.22 61.19 59.48
CoOp[28] 56.75 60.17 64.13 62.66 60.14
DPOD (Ours) 58.92 61.12 65.07 62.90 62.70
Evaluation on standard setting for all domains. Our proposed DPOD is a generalized framework, where we train our model with data from all the diverse domains available. Along with evaluating on all the domains combined, the framework can also generalize to domains of interest, by leveraging information of other domains. First, we evaluate our model on image-text pairs available from all the domains combined, and report the results in Table 1. Here, we assume that we know the domain label of the test data. The experiment is run $5$ times with random sampling of training data, and the mean performance is reported. We compare our proposed framework with the previous SOTA approaches on NewsCLIPpings dataset, namely, CLIP-FT [1], SSDL [16], DT-Transformer [17], MNSL[25], DCD[10], NewsCLIPpings[15]. For CLIP-FT, the entire CLIP model is fine-tuned along with two additional layers with the available training data. This is the performance without using the open-domain data. For all the other approaches, we directly report the accuracies reported in their papers. For the scenario where domain knowledge of the test sample is unknown, our model gives an accuracy of $69.13\%$, which is comparable to the performance achieved when the test domain information is known.
For fair comparison, we also perform our experiments with different variants of the CLIP model, namely, ViT-B/32, ViT-B/16 and RN-50. We observe that the proposed framework consistently outperforms the SOTA by a significant margin when similar backbones are used for comparison, in both the settings.
Figure 2: (a) Computation of the joint embedding of the images and text for a particular domain. These embedding are used in Stage 2 and 3 of our DPOD framework. (b) Similarity of the learnt domain-specific prompts of different domains. We observe that the sets Football, Sport, Sports and Healthcare Network, Healthcare Medicine have similar domain-specific prompts among themselves, but they have low similarity with each other and also with Music.
Evaluation on a specific domain. Now, we consider the case where an individual might be interested in news from a specific domain. We report in Table 2 how well our model utilizes out-of-domain samples in order to predict the veracity of the news belonging to a desired domain. To highlight the problem of data imbalance across different domains, we consider $5$ domains with the most datapoints, namely, Sport, Law-crime, International Relations, World, Arts&Culture, along with $5$ domains with the least data samples, namely, Fashion. TV-Radio, Education, Money, Travel. In the absence of comparative approaches, we experiment with the following strong baselines:
1) Clip-FT [1]: without the open-domain information for fair comparison with the other approaches.
2) CoOp [28]: First, the prompts $\{V_{1},V_{2},V_{3}\}$ are learnt using training data from all the domains. Then, these prompts are frozen for evaluation on the target domains. For fair comparison with our approach, CoOp is also used with A-CLIP. All the models here are trained on the entire dataset. From Table 2, we observe that DPOD consistently outperforms the other compared methods for all the individual domains, irrespective of the number of training examples available for the particular domain, highlighting the effectiveness of the proposed domain-specific prompt learning framework.
Table 3: Comparison of accuracies (in %) where we train the model on 90% domains and evaluate its performance on the rest 10% unseen domains.
Methods Politics Law Crime Arts & Cul. Sports Music
CLIP-FT 68.31 61.65 64.87 71.05 63.33
DPOD 72.29 64.63 71.48 73.18 66.67
Figure 3: Qualitative results of the proposed DPOD Model. First three columns on the left with green border are examples of successful predictions, and the three on the right with pink border are examples of failure cases.
Figure 4: The yellow regions show the portion of the image and text where A-CLIP attends to during classification.
Generalization to unseen domains. Till now, we have assumed that the test domain is already present during training. Here, we evaluate the performance of our model when presented with completely new domains not included in the training set. For this, we train our model using samples from 90% of the domains, and test it on the remaining 10% unseen domains, using the most similar domain-specific prompt (based on cosine similarity). For a fair comparison, we applied a similar setting for CLIP-FT, and report the results in Table 3. Our proposed DPOD consistently outperforms the state-of-the-art CLIP-FT for all the unseen domain categories.
Handling Domain-Label Inconsistencies. Often, there are inconsistencies in the domain labels, which may be due to crowd-sourcing for obtaining labels. For example, in the NewsCLIPpings dataset, there are domain labels like football, sports, sport and also Healthcare Network, Healthcare Medicine, etc. which should ideally have the same domain labels. Fig. 1(b) shows the cosine-similarity matrix that is calculated by taking the cosine similarity between the learnt domain-specific prompts of the corresponding two domains. We observe that the learnt domain-specific prompts inherently captures the semantic information of the domains, thereby bringing the similar domains closer, thus helping in their training. Our proposed DPOD does not require the exact domain names, and thus can work even if it is difficult to specify the domains, and dummy domain labels (like $D_{1},D_{2},etc.$) are used instead.
Ablation Studies. Here, we analyze the importance of Stage 1 (label-aware alignment) and Stage 2 (semantic domain vectors) for the all domain scenario in Table 5. We observe that removing either of the stages deteriorates the overall performance. The second row in this table is generated by initializing $\{{V}_{1},{V}_{2},{V}_{3}\}$ with “A photo of” and appending it with a fourth vector ${V}_{4}$, which is obtained by passing a unique one-hot vector corresponding to each training domains. This also implies that if the domains are treated independently, then they cannot fully benefit from the training examples of closely related domains as in the proposed DPOD. The importance of domain-specific prompts can also be observed from Table 5, which shows the performance of prompt $\{{V}_{1},{V}_{2},{V}_{3}\}$ and $\{{V}_{1},{V}_{2},{V}_{3},{V}_{4}\}$, where ${V}_{4}$ is another learnable prompt, but not a domain specific one as in DPOD (third row). Thus, all the proposed modules contribute to the improved performance of DPOD framework.
Table 4: Usefulness of the two stages of the DPOD framework. Effectiveness of Stage 2 was evaluated by taking unique one-hot vectors corresponding to the unique domains.
Table 5: Usefulness of domain-specific prompt tuning. Here the models have been trained on data from all the domains.
Stage 1 Stage 2 Accuracy(%)
$\times$ 68.205
$\times$ 68.923
70.811
Prompts Test
Politics Sport All Domain
{${V}_{1}$, ${V}_{2}$, ${V}_{3}$} 77.027 69.68 69.349
{${V}_{1}$, ${V}_{2}$, ${V}_{3}$, ${V}_{4}$} 75 68.617 70.351
DPOD: {${V}_{1}$, ${V}_{2}$, ${V}_{3}$, ${V}_{D}$} 78.378 71.011 70.811
Table 5: Usefulness of domain-specific prompt tuning. Here the models have been trained on data from all the domains.
Qualitative Results and Limitations. Here, we provide some qualitative results of our framework. Fig. 3 (left with green border) shows few examples of correct predictions. We observe that the model demonstrates the capability to accurately assess the veracity of a news item even in complex situations, as in the first example of a promotional event for a different brand. Fig. 3 (right with pink border) shows few examples of incorrect predictions by our model. We observe that in some cases, the model struggles when multiple elements mentioned in the caption are present in the image, as Koala in the second failure case. In Fig. 4, we present a few qualitative results illustrating where A-CLIP focuses on in both the image and text after Stage 1 in the proposed DPOD method. We observe that, in most cases, A-CLIP attends to regions (marked in yellow) that play a crucial role in decision making.
6 Conclusion
Here, we proposed a novel framework for the challenging MFND task. First, label-aware alignment of the data is achieved using the widely-popular CLIP model to obtain generalized features. Further, we propose to learn generic as well as domain-specific prompts to classify the input image-text pairs. Experiments show that the proposed DPOD achieves the new state-of-the-art for this challenging socially relevant MFND task.